OpenStudy (ramennoodle01):

A sailboat costs $35,000.You pay 20% down and amortize the rest with equal monthly payments over a 12-year period. If you must pay 8.75% compounded monthly, what is your monthly payment? How much interest will you pay?

OpenStudy (mathmate):

20% down means 35000*0.2=7000 down. What's left to be amortized is P=35000-7000=28000 (present value) Let A=monthly payment, n=number of months (144) i=monthly interest=8.75%/12=0.0875/12 R=monthly compounding factor = 1+i Then \(A=P(R^n)i/(R^n-1)\) Take out your calculator and find A.